### 06 Dec 2023 v1

# On the geometry of the free factor graph for ${\rm{Aut}}(F_{N})$

Abstract
Let $\Phi$ be a pseudo-Anosov diffeomorphism of a compact (possibily non-orientable) surface $\Sigma$ with one boundary component. We show that if $b\in\pi_{1}(\Sigma)$ is the boundary word, $\phi\in{\rm{Aut}}(\pi_{1}(\Sigma))$ is a representative of $\Phi$ fixing $b$, and ${\rm{ad}}_{b}$ denotes conjugation by $b$, then the orbits of $\langle\phi,{\rm{ad}}_{b}\rangle\cong\mathbb{Z}^{2}$ in the graph of free factors of $\pi_{1}(\Sigma)$ are quasi-isometrically embedded. It follows that for $N\geq 2$ the free factor graph for ${\rm{Aut}}(F_{N})$ is not hyperbolic, in contrast to the ${\rm{Out}}(F_{N})$ case.
1 Introduction
The set of nontrivial proper free factors of a free group $F_{N}$ can be ordered by inclusion and the geometric realisation of the resulting poset is the free factor complex for ${\rm{Aut}}(F_{N})$. This complex was introduced by Hatcher and Vogtmann [HV98] who proved, in analogy with the Solomon–Tits theorem for the Tits building associated to ${\rm{GL}}(N,\mathbb{Z})$, that this complex has the homotopy type of a wedge of spheres of dimension $N-2$. Our focus here is on the large-scale geometry of the complex rather than its topology. This geometry is captured entirely by the 1-skeleton, i.e. the free factor graph $\mathcal{AF}_{N}$, metrized as a length space with edges of length $1$. Thus the vertices of $\mathcal{AF}_{N}$ are the nontrivial, proper free factors of $F_{N}$ and there is an edge joining $A$ to $B$ if $A<B$ or $B<A$. (When $N=2$ this definition is modified – see Section 2.)
There is a natural action of ${\rm{Aut}}(F_{N})$ on $\mathcal{AF}_{N}$. The quotient of $\mathcal{AF}_{N}$ by the subgroup of inner automorphisms is called the free factor graph for ${\rm{Out}}(F_{N})$. This graph, which is denoted by $\mathcal{OF}_{N}$, has emerged in recent years as a pivotal object in the study of ${\rm{Out}}(F_{N})$. Much of its importance derives from the following fundamental result of Bestvina and Feighn [BF14].
Theorem A (Bestvina–Feighn [BF14]).
The free factor graph $\mathcal{OF}_{N}$ is Gromov-hyperbolic. The fully irreducible elements of ${\rm{Out}}(F_{N})$ act as loxodromic isometries of $\mathcal{OF}_{N}$ (i.e. have quasi-isometrically embedded orbits) while every other element has a finite orbit.
The natural map ${\rm{Out}}(F_{N})\to{\rm{Isom}}(\mathcal{OF}_{N})$ is an isomorphism [BB23], so Theorem A describes all possible actions of cyclic groups on $\mathcal{OF}_{N}$. As the graph is hyperbolic, we also know that $\mathbb{Z}^{2}$ cannot with act quasi-isometric orbits. In fact, results in the literature [BB23, HM20] tell us something much stronger:
Theorem B.
If $r\geq 2$, then every faithful action of $\mathbb{Z}^{r}$ by isometries on $\mathcal{OF}_{N}$ has a finite orbit.
Proof.
The natural map ${\rm{Out}}(F_{N})\to{\rm{Isom}}(\mathcal{OF}_{N})$ is an isomorphism [BB23], so it suffices to prove this theorem for abelian subgroups of ${\rm{Out}}(F_{N})$. The centralisers in ${\rm{Out}}(F_{N})$ of fully irreducible elements are virtually cyclic, so an abelian subgroup of rank greater than $1$ contains no fully irreducible elements. Handel and Mosher [HM20] proved that if a finitely generated subgroup $G<{\rm{Out}}(F_{N})$ contains no fully irreducible elements, then $G$ has a finite orbit in $\mathcal{OF}_{N}$ (a fact that remains true even if $G$ is not finitely generated [Hor16]). ∎
The purpose of this article is to prove results concerning $\mathcal{AF}_{N}$ that contrast sharply with Theorems A and B. We shall construct actions of $\mathbb{Z}^{2}$ on $\mathcal{AF}_{N}$ such that the orbits are free and metrically undistorted. We shall also establish a criterion that tells us certain inner automorphisms act as isometries with quasi-isometrically embedded orbits, just as all fully irreducible automorphisms do. For the moment, however, we are unable to offer a concise classification of all the isometries that have such orbits. (The natural map ${\rm{Aut}}(F_{N})\to{\rm{Isom}}(\mathcal{AF}_{N})$ is an isomorphism so, as in the ${\rm{Out}}$ case, there ought to be a classification in purely algebraic terms.)
The abelian subgroups $\mathbb{Z}^{2}\hookrightarrow{\rm{Aut}}(F_{N})$ that we shall focus on are constructed using pseudo-Anosov diffeomorphisms of compact surfaces with one boundary component. Let $\Sigma$ be such a surface with Euler characteristic $1-N$. Let $\text{{\rm{Mod}}}(\Sigma)$ be the mapping class group of $\Sigma$, that is, $\pi_{0}$ of the subgroup of ${\rm{Diff}}(\Sigma)$ that fixes $\partial\Sigma$ pointwise. Then by fixing a basepoint on $\partial\Sigma$ we obtain an identification $\pi_{1}\Sigma=F_{N}$ and a monomorphism $\text{{\rm{Mod}}}(\Sigma)\to{\rm{Aut}}(F_{N})$. The Dehn twist in the boundary of $\Sigma$ is central in $\text{{\rm{Mod}}}(\Sigma)$. We are interested in the $\mathbb{Z}^{2}$ subgroup that this twist generates with any pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$. The Dehn twist acts on $\pi_{1}\Sigma=F_{N}$ as an inner automorphism ${\rm{ad}}_{b}$, where $b\in F_{N}$ is the boundary loop.
Theorem C.
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
Corollary D.
For $N\geq 2$, the free factor graph $\mathcal{AF}_{N}$ is not Gromov-hyperbolic.
These results should be compared with the work of Hamenstädt [Ham21], who constructed similar quasi-flats in spotted disc complexes of handlebodies and in certain sphere complexes.
A crucial property of the boundary element $b\in F_{N}$ is that it is filling in the sense that it is not contained in any proper free factor of $F_{N}$. Given any filling element $b$, we analyse the action of ${\rm{ad}}_{b}$ on $\mathcal{AF}_{N}$ with the aim or showing that the orbits of $\langle{\rm{ad}}_{b}\rangle$ give quasi-isometric embeddings of $\mathbb{Z}$. The key tool in this analysis is the notion of $b$-reduced decomposition for elements $w\in F_{N}$, which we introduce in Section 3. Using these decompositions, we define an integer-valued invariant $[A]_{b}$ of free factors $A$ and we use this to define a Lipschitz retraction of $\mathcal{AF}_{N}$ onto the $\langle{\rm{ad}}_{b}\rangle$-orbit of a vertex $V$; roughly speaking, this retraction sends the vertex $A$ to ${\rm{ad}}_{b}^{n}(V)$ if $n=[A]_{b}$.
Theorem E.
If an element $b\in F_{N}$ is not contained in any proper free factor, then there is a Lipschitz retraction of $\mathcal{AF}_{N}$ onto each orbit of $\langle{\rm{ad}}_{b}\rangle$. In particular, these orbits are quasi-isometrically embedded.
In Section 2 we recall classical results of Whitehead and others describing the structure of elements $b\in F_{N}$ that are not contained in any proper free factor. This structure controls the behaviour of the function $A\mapsto[A]_{b}$ that is used in Section 3 to prove Theorem E. With Theorem E in hand, one feels that Theorem C ought to follow easily, since orbits of the fully irreducible automorphism $\phi$ project to quasi-geodesics in $\mathcal{OF}_{N}$, but the details are slightly delicate; this is explained in Section 4.
The fact that we cannot offer a concise classification of the isometries of $\mathcal{AF}_{N}$ is symptomatic of the fact that the large-scale geometry of $\mathcal{AF}_{N}$ is poorly understood. In Section 5 we highlight some of the begging questions in this regard.
Acknowledgements. This paper is dedicated to Slava Grigorchuk on his 70th birthday, with thanks for his many highly-original contributions to group theory. The first author acknowledges the support of the National Science Foundation under the grant number DMS-2304774. The third author is supported by the Royal Society of Great Britain through a University Research Fellowship.
2 Background
We assume that the reader is familiar with the rudiments of the theory of free-group automorphisms.
We write $F_{N}$ to denote the free group on $N$ generators. Let $\mathcal{S}$ be a basis of $F_{N}$. Every element $w\in F_{N}$ is represented by a unique reduced word in the letters $\mathcal{S}\cup\mathcal{S}^{-1}$; the length of this word is denoted by $|w|_{\mathcal{S}}$. We write $=$ to denote equality in $F_{N}$ and $\equiv$ to denote equality as words. A subgroup $A\leq F_{N}$ is a free factor if it is generated by a subset of a basis, and is a proper free factor if $A\neq F_{N}$.
The natural definition of $\mathcal{AF}_{N}$ that we gave in the introduction is unsatisfactory in the case $N=2$, since there are no edges. We remedy this by decreeing that a pair of vertices $\langle u\rangle$ and $\langle v\rangle$ are to be connected by an edge if and only if $\{u,v\}$ is a basis for $F_{2}$. Then $\mathcal{OF}_{2}$ is defined to be the quotient by the action of the inner automorphisms. With these conventions, $\mathcal{OF}_{2}$ is the Farey graph with the standard action of ${\rm{Out}}(F_{2})\cong{\rm{GL}}(2,\mathbb{Z})$.
2.1 Whitehead graphs, word length, and the cut-vertex lemma
An element of $F_{N}$ is primitive if it belongs to some basis, and is simple if it belongs to some proper free factor. Every element $w$ has a cyclic reduction $w_{0}$, obtained by writing $w\equiv w_{1}w_{0}w_{1}^{-1}$ with $w_{1}$ as long as possible. An element is cyclically reduced (with respect to the basis $\mathcal{S}$) if $w=w_{0}$. We use ${\mathrm{Wh}}_{\mathcal{S}}(w)$ to denote the cyclic Whitehead graph of $w$. This is the graph with vertex set $\mathcal{S}\cup\mathcal{S}^{-1}$ that has an edge from $x$ to $y^{-1}$ if $xy$ is a subword of the cyclic reduction $w_{0}$ or if $x$ is the last letter of $w_{0}$ and $y$ is the first letter of $w_{0}$. A vertex $v$ of a graph is a cut vertex if the full subgraph spanned by vertices not equal to $v$ is disconnected. The famous lemma below was stated by Whitehead for primitive elements [Whi36] and the generalization to simple elements was observed by various authors (see, e.g. [Mar95, Proposition 49] or [Ota92, Sta99, HW19]). (For readers who pursue the references: Whitehead uses the word simple differently to modern authors—his simple sets are subsets of bases.)
Lemma 2.1 (Whitehead’s Cut-Vertex Lemma [Whi36]).
Let $N\geq 2$. If $w$ is a simple element of $F_{N}$ then for any basis $\mathcal{S}$ of $F_{N}$ the cyclic Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(w)$ contains a cut vertex.
If the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(w)$ contains a cut vertex that is not isolated, then there is a Whitehead automorphism that reduces the cyclic length of $w$. Using this, one can prove the following standard proposition, which can be seen as a partial converse to Whitehead’s lemma.
Proposition 2.2.
Let $N\geq 2$. If $w\in F_{N}$ is not contained in a proper free factor and $|w|_{\mathcal{S}}\leq|\phi(w)|_{\mathcal{S}}$ for all $\phi\in{\rm{Aut}}(F_{N})$, then ${\mathrm{Wh}}_{\mathcal{S}}(w)$ contains no cut vertex.
Note that any graph with at least 3 vertices that is disconnected contains a cut vertex. As we are working with $N\geq 2$, all of the Whitehead graphs we consider have at least 4 vertices, and so those without cut vertices are connected.
3 Orbits in $\mathcal{AF}_{N}$ of filling inner automorphisms
The goal of this section is to prove Theorem E: if $b\in F_{N}$ is not contained in a proper free factor, then there is a Lipschtiz retraction of $\mathcal{AF}_{N}$ onto each $\langle{\rm{ad}}_{b}\rangle$-orbit. These retractions will be constructed using $b$-reduced decompositions.
3.1 $b$-reduced decompositions and a cancellation lemma
Definition 3.1 ($b$-reduced decomposition).
Fix a basis $\mathcal{S}$ of $F_{N}$. Given a cyclically reduced word $b$ and a reduced word $w$, representing elements of $F_{N}$, the $b$-reduced decomposition of $w$ (with respect to $\mathcal{S}$) is the decomposition
$w\equiv b^{k}w_{b}b^{-k}$
(equality as reduced words) such that $|k|$ is maximal. Define $[w]^{\mathcal{S}}_{b}:=k\in\mathbb{Z}$. Given a subgroup $A\subset F_{N}$, define
$[A]^{\mathcal{S}}_{b}:=\sup\{[a]^{\mathcal{S}}_{b}:a\in A\},$
allowing $[A]^{\mathcal{S}}_{b}=\infty$.
We shall drop the superscript $\mathcal{S}$ from $[w]^{\mathcal{S}}_{b}$ and $[A]_{b}^{\mathcal{S}}$ when there is no danger of ambiguity.
Remark 3.2 (Geometric Interpretation).
Consider the action of $F_{N}$ on the Cayley tree associated to $\mathcal{S}$. As $b$ is cyclically reduced, it acts as translation through a distance $|b|$ in an axis $X_{b}$ that passes through the basepoint $1$ and contains each vertex $b^{i}$. Roughly speaking, $[a]_{b}$ records (with sign) the time that the path from $1$ to the axis $X_{a}$ of $a$ spends on $X_{b}$. More precisely, the orthogonal projection of $X_{a}$ onto $X_{b}$ is contained in a minimal interval of the form $[b^{i},b^{j}]$ and $[a]_{b}$ is $i$ if $i>0$, is $j$ if $j<0$, and is 00 if $i\leq 0\leq j$ – see the proof of Lemma 3.3. Note that the projection of $X_{a}$ to $X_{b}$ is a point if $X_{a}\cap X_{b}=\emptyset$ and equals $X_{a}\cap X_{b}$ otherwise.
The minimal invariant subtree $T_{A}$ for a subgroup $A\leq F_{N}$ is the union of the axes $X_{a}$ with $a\in A$. If $T_{A}\cap X_{b}$ is compact, as it will be if $A$ is finitely generated and $A\cap\,\langle b\rangle=1$, then either it is empty, in which case the projection of $T_{A}$ is a point and $[A]_{b}=[a]_{b}$ for all $a\in A$, or else $T_{A}\cap X_{b}$ is contained in a minimal interval of the form $[b^{I},b^{J}]$ and from the geometric description of $[a]_{b}$ we have $I\leq[A]_{b}\leq J$. These observations lead to the philosophically important approximation
[…]
where $\pi_{X_{b}}$ denotes orthogonal projection to $X_{b}$ and the constants implicit in the approximation depend on the length of $T_{A}\cap X_{b}$.
[…]
Figure 1: The proof of the cancellation lemma.
Lemma 3.3 (Cancellation Lemma).
Suppose that $b\in F_{N}$ is not contained in a proper free factor and let $\mathcal{S}$ be a basis of $F_{N}$ that minimizes $|b|_{\mathcal{S}}$. Suppose that $a\in F_{N}$ is contained in a proper free factor and $[a]_{b}^{\mathcal{S}}=0$. Then the first and last $|b|_{\mathcal{S}}+1$ letters of $w\equiv b^{3}\cdot a\cdot b^{-3}$ remain after reducing $w$. In particular, $[b^{3}ab^{-3}]_{b}^{\mathcal{S}}\geq 1$.
Proof.
Let $T$ be the Cayley tree for $F_{N}$ with respect to the basis $\mathcal{S}$. As $b$ is cyclically reduced, the axis $X_{b}$ of $b$ passes through the identity vertex $1\in T$. Let $X_{a}$ be the axis for $a$. Let $p$ be the shortest path from $1$ to $X_{a}$. The intersection of $p$ with $X_{b}$ is of length at most $|b|_{\mathcal{S}}-1$, as the reduced word representing $a$ starts with the word labelling $p$ and ends with its inverse and $[a]_{b}=0$. Furthermore, the intersection of $X_{a}$ with $X_{b}$ can have length at most $|b|_{\mathcal{S}}$: to see this, note that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ does not contain a cut vertex by Proposition 2.2, whereas ${\mathrm{Wh}}_{\mathcal{S}}(a)$ does by the Cut Vertex Lemma, so $X_{a}$ cannot contain all the turns (pairs of consecutive of labels from $\mathcal{S}\cup\mathcal{S}^{-1}$) that appear along $X_{b}$.
Let $y$ by the furthest point along $X_{b}$ in the $b^{-1}$ direction that is contained in $p\cup X_{a}$. By the above, $d(1,y)\leq 2|b|_{\mathcal{S}}-1$. It follows that the distance from $1$ to $b^{3}y$ (in the positive direction along $X_{b}$) is at least $|b|_{\mathcal{S}}+1$. This implies that the path from $1$ to the axis of $b^{3}ab^{-3}$ contains the interval $[1,b^{3}y]$. Again, as the path from 1 to the axis gives the initial and terminal (in reverse) subwords of $b^{3}ab^{-3}$, the result follows. ∎
Remark 3.4.
Consider the orbit $\mathcal{O}=\{\langle b^{k}ab^{-k}\rangle:k\in\mathbb{Z}\}$ of a free factor $\langle a\rangle$ generated by a primitive element under the action of ${\rm{ad}}_{b}$. It is not hard to see that the map $f:\mathcal{O}\to\mathbb{Z}$ given by $f(w)=[w]_{b}$ is injective outside of $f^{-1}(0)$. And under the assumptions of the above lemma, $|f^{-1}(0)|\leq 3$.
Proposition 3.5.
Suppose $b\in F_{N}$ is cyclically reduced with respect to ${\mathcal{S}}$ and that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ is connected with no cut vertex. Let $T$ be the Cayley tree with respect to $\mathcal{S}$ and let $X_{b}$ be the axis of $b$ in $T$. If $A\leq F_{N}$ is a proper free factor and $T_{A}$ is the minimal subtree of $A$ in $T$, then
  1. 1.
    $|T_{A}\cap X_{b}|\leq|b|_{{\mathcal{S}}}$.
  2. 2.
    For any two elements $a_{1},a_{2}\in A$,
    $|[a_{1}]_{b}^{\mathcal{S}}-[a_{2}]_{b}^{\mathcal{S}}|\leq 1.$
In particular, $[A]_{b}^{\mathcal{S}}\in\mathbb{Z}$.
Proof.
If $T_{A}\cap X_{b}=\emptyset$ then $X_{a}$ has the same projection to $X_{b}$ for all $a\in A$. In this case part (1) is trivial and $[a_{1}]_{b}^{\mathcal{S}}=[a_{2}]_{b}^{\mathcal{S}}$ for all $a_{1},a_{2},\in A$ by the Geometric Interpretation (Remark 3.2). Otherwise, the projection $T_{A}$ to $X_{b}$ is a compact interval equal to $T_{A}\cap X_{b}$. As $T_{A}\cap X_{b}$ is compact, there exists $a\in A$ such that the axis of $a$ contains $T_{A}\cap X_{b}$ by [Pau89, Lemma 4.3]. As every element of $A$ is simple, as in the preceding proof this would contradict the Cut Vertex Lemma if $|T_{A}\cap X_{b}|>|b|_{{\mathcal{S}}}$. This proves (1). For (2), it then suffices to show that if $a_{1},a_{2}\in F_{N}$ are elements of $F_{N}$ and the projections of $X_{a_{1}}$ and $X_{a_{2}}$ to $X_{b}$ are contained in an interval of length at most $|b|_{{\mathcal{S}}}$, then $|[a_{1}]_{b}^{\mathcal{S}}-[a_{2}]_{b}^{\mathcal{S}}|\leq 1$. This again follows from the Geometric Interpretation: $T_{A}\cap X_{b}$ is contained in $(b^{i-1},b^{i+1}]$ for some $i<0$, or contained in $[b^{i-1},b^{i+1})$ for some $i\geq 0$. In the first case, one can check that $[a]_{b}^{\mathcal{S}}\in\{i,i+1\}$ for all $a\in A$, and in the second $[a]_{b}^{\mathcal{S}}\in\{i-1,i\}$ for all $a\in A$. ∎
Theorem 3.6.
Suppose that $b\in F_{N}$ is not contained in a proper free factor and that $\mathcal{S}$ is a basis that minimizes $|b|_{\mathcal{S}}$. Then the map $\mathcal{AF}\to\mathbb{Z}$ given by $A\mapsto[A]^{\mathcal{S}}_{b}$ is 1–Lipschitz if $N\geq 3$ and is 2–Lipschitz if $N=2$.
Proof.
Proposition 2.2 tells us that Proposition 3.5 applies to $b$ and Proposition 3.5(2) assures us that if $A<B$ then $|[A]^{\mathcal{S}}_{b}-[B]^{\mathcal{S}}_{b}|\leq 1$; this covers the case $N=3$. For the case $N=2$, a different argument is needed, because $\mathcal{AF}_{N}$ is defined differently. In this case, the vertex $\langle a_{1}\rangle$ is adjacent to $\langle a_{2}\rangle$ if $\{a_{1},a_{2}\}$ is a basis. As in Proposition 3.5(1), we know that the projection $\pi_{X_{b}}(X_{a_{i}})$ of each axis to $X_{b}$ has length at most $|b|_{\mathcal{S}}$. We claim that furthermore, if the projections are disjoint then the arc connecting $\pi_{X_{b}}(X_{a_{1}})$ to $\pi_{X_{b}}(X_{a_{2}})$ is length at most $|b|_{\mathcal{S}}$. This claim clearly gives a uniform bound on the difference between $[a_{1}]_{b}$ and $[a_{2}]_{b}$, and with a similar case-by-case analysis as in Proposition 3.5 one can check that $[a_{1}]_{b}$ and $[a_{2}]_{b}$ differ by at most $2$.
To prove the claim, if $X_{a_{1}}$ and $X_{a_{2}}$ do not intersect, then the bridge between them is contained in the axis for $a_{1}a_{2}$ (see [CM87] Figure 4). It follows that if $\pi_{X_{b}}(X_{a_{1}})$ and $\pi_{X_{b}}(X_{a_{2}})$ do not intersect, then the arc between the projections is contained in $\pi_{X_{b}}(X_{a_{1}a_{2}})$. Since $a_{1}a_{2}$ is primitive, Proposition 3.5(1) tells us that $\pi_{X_{b}}(X_{a_{1}a_{2}})$ has length at most $|b|_{\mathcal{S}}$, and we have proved the claim.∎
3.2 Change of Basis
Proposition 3.7.
Let $b\in F_{N}$ be an element that is not contained in a proper free factor. Let $\mathcal{S}$ and $\mathcal{T}$ be bases of $F_{N}$ which minimize the word length of $b$. Then there exists a constant $K$ such that for all $A\in\mathcal{AF}_{N}$ we have:
$[A]_{b}^{\mathcal{T}}-K\leq[A]_{b}^{\mathcal{S}}\leq[A]_{b}^{\mathcal{T}}+K.$
Proof.
Let $T^{\mathcal{S}}$ and $T^{\mathcal{T}}$ be the Cayley trees of $F_{N}$ given by $\mathcal{S}$ and $\mathcal{T}$, respectively. Let $f:T^{\mathcal{S}}\to T^{\mathcal{T}}$ be the map that sends each edge of $T^{\mathcal{S}}$ to the geodesic in $T^{\mathcal{T}}$ with the same endpoints. If $p=[1,g]_{\mathcal{S}}$ is a path in $T^{\mathcal{S}}$ from 1 to the axis of an element $a\in F_{N}$, then the bounded backtracking property (see [Coo87, GJLL98]) implies that $g$ lies at distance at most $C$ from the axis of $a$ in $T^{\mathcal{T}}$, where $C$ is the bounded backtracking constant. Therefore, as $|b|_{\mathcal{S}}=|b|_{\mathcal{T}}$, the proposition reduces to showing that for any path $p$ based at 1, the reduction of $f(p)$ spends approximately the same amount of time travelling along the axis of $b$ in $T^{\mathcal{T}}$ as the path $p$ does along the axis of $b$ in $T^{\mathcal{S}}$ (in the same direction). This is reasonably standard: decompose such a path $p$ as $p=p_{1}p_{2}p_{3}$, where $p_{1}=[1,b^{k}]_{\mathcal{S}}$ for some $k$, the path $p_{2}$ is a segment of length $\leq|b|_{\mathcal{S}}$, and $p_{3}$ is a path disjoint from the axis $X^{\mathcal{S}}_{b}\subset T^{\mathcal{S}}$ except for its initial endpoint. As $f$ is bi-Lipschitz on the vertex set, $p_{3}$ can further be decomposed as $p_{3}^{\prime}p_{3}^{\prime\prime}$, where $f(p_{3}^{\prime\prime})$ is disjoint from $X^{\mathcal{T}}_{b}$ and $p_{3}^{\prime}$ is of uniformly bounded length. Then the intersection of the reduction of $f(p)$ with $X^{\mathcal{T}}_{b}$ in $T_{\mathcal{T}}$ is $[1,b^{k}]_{\mathcal{T}}$, up to cancellation/concatenation with a subpath of $f(p_{2}p_{3}^{\prime})$, which is of uniformly bounded length. ∎
4 Building quasi-flats
In this section we will build quasi-flats in $\mathcal{AF}_{N}$. The first proposition is quite technical: exceptionally, in the proof of this proposition we will assume the reader is familiar with the stable train tracks of [BH92].
Proposition 4.1.
Let $\Sigma$ be a compact surface with one boundary component, and fix a basepoint on the boundary giving an identification $\pi_{1}\Sigma=F_{N}$. Let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi_{0}\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, and let $[\phi_{0}]$ be its outer automorphism class. Then, there exists a basis $\mathcal{S}$ of $F_{N}$ and a representative $\phi$ of $[\phi_{0}]$ fixing $b$, such that:
  1. 1.
    The basis $\mathcal{S}$ minimizes $|b|_{\mathcal{S}}$.
  2. 2.
    For all $A\in\mathcal{AF}_{N}$ there exists a constant $M_{A}$ such that
    $|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq M_{A}$
    for all $k,r\in\mathbb{Z}$.
  3. 3.
    $\phi$ is the unique representative of $[\phi_{0}]$ that fixes $b$ and acts on the Gromov boundary $\partial F_{N}$ so that the endpoints of the infinite words $b^{\infty}$ and $b^{-\infty}$ are both non-attracting fixed points.
Proof.
In Sections 1-3 of [BH92], it is shown that an automorphism $\phi_{0}$ as above has a stable train track representative $f\colon G\to G$ with exactly one indivisible Nielsen path (iNP) $\rho$ which, as the automorphism is geometric, forms a loop representing $b$ based at a point $x\in G$ that is fixed by $f$. Furthermore, $\rho$ crosses every edge in $G$ exactly twice (see [BH92, Lemma 3.9] and the discussion in [BH92, Section 4]). Identifying $\pi_{1}(G,x)$ with $F_{N}$, we define $\phi$ to be the automorphism induced by $f_{*}\colon\pi_{1}(G,x)\to\pi_{1}(G,x)$ and take $\mathcal{S}$ to be a basis of $F_{N}$ determined by a choice of maximal tree $T$ in $G$ together with an orientation of the edges in $G\smallsetminus T$. Each edge-loop based at $x$ determines a word in the letters ${\mathcal{S}}^{\pm 1}$ whose length is the number of edges of $G\smallsetminus T$ crossed by the loop. In particular, $|b|_{\mathcal{S}}=2N$. Since $b$ is the boundary for a surface of Euler characteristic $1-N$, this tells us that $\mathcal{S}$ minimizes the length of $b$.
We now prove the second part with $r\geq 0$. There exists a legal loop $l$ in $G$ based at $x$ that crosses some edge exactly once (this is a variant of Francaviglia and Martino’s Sausage Lemma [FM11, Lemma 3.14]). Any such loop represents a primitive element in $F_{N}$. As in Section 2 of [BH92], we lift $f$ to a map $\tilde{f}\colon\tilde{G}\to\tilde{G}$ on the universal cover that represents $\phi$ (i.e. $\tilde{f}(gy)=\phi(g)\tilde{f}(y)$ for all $y\in\tilde{G}$). The map $\tilde{f}$ fixes a lift $\tilde{x}$ of our basepoint $x\in G$, and the axis of $b$ (acting as a deck transformation of $\tilde{G}$) contains $\tilde{x}$. As the loop $l$ is based at $x$ its axis $\operatorname{Axis}_{\tilde{G}}(l)$ in $\tilde{G}$ also contains $\tilde{x}$. As $l$ is legal and $\tilde{x}$ is fixed by $\tilde{f}$, the axis of $\phi^{r}(l)$ also contains $\tilde{x}$ for all $r\geq 0$. It follows that iterated images of $l$ are cyclically reduced, so that $[\phi^{r}(l)]_{b}=0$ for all $r\geq 0$. For $k\in\mathbb{Z}$, the axis of ${\rm{ad}}_{b}^{k}(l)$ is also legal and contains the point $b^{k}\tilde{x}$, so that $k-1\leq[{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$. As $b^{k}\tilde{x}$ is also fixed under $\tilde{f}$, this implies that $k-1\leq[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$ for all $k\in\mathbb{Z}$ and $r\geq 0$. As the map $\mathcal{AF}_{N}\to\mathbb{Z}$ induced by $[A]\mapsto[A]_{b}$ is 2-Lipschitz, this implies that for all $A\in\mathcal{AF}_{N}$ there exists $C_{A}$ such that $|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}$ for all $r\geq 0$ and $k\in\mathbb{Z}$.
Note that any possible choices for representatives of $[\phi_{0}]$ fixing $b$ differ by a power of ${\rm{ad}}_{b}$. If $\phi$ is the representative chosen above then as $\tilde{f}(b^{n}\tilde{x})=b^{n}\tilde{x}$ for all $n\in\mathbb{Z}$, both ends of the axis of $b$ in $\tilde{G}$ are non-attracting under the action of $\phi$ on $\partial F_{N}$. For $k\neq 0$, one end of the axis will be attracting and the other will be repelling under the action of ${\rm{ad}}_{b}^{k}\phi$ on the boundary. This establishes (3).
To deal with the case where $r<0$ in (2), we run the entire argument again for $[\phi_{0}]^{-1}$. This provides a representative $\psi$ of $[\phi_{0}]^{-1}$ fixing $b$, such that for all $A\in\mathcal{AF}_{N}$ there exists a constant $C_{A}^{\prime}$ such that $|[\psi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}^{\prime}$ for all $k\in\mathbb{Z}$ and $r\geq 0$. Note that although you might be working with a different basis $\mathcal{S}^{\prime}$ on which $b$ is minimal length, the projections $\mathcal{AF}_{N}\to\mathbb{Z}$ are coarsely equivalent by Proposition 3.7. Furthermore, this is the unique representative of $[\phi^{\prime}]^{-1}$ fixing $b$ such that $b^{\infty}$ and $b^{-\infty}$ are nonattracting under $\partial\psi$. This implies that $\psi=\phi^{-1}$, so we can finish part (2) by taking $M_{A}:=\max\{C_{A},C_{A}^{\prime}\}$.
We finally have all of the tools that we need to prove Theorem C, which we restate for the reader’s convenience.
Theorem 4.2.
Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$. Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
Proof.
Note that it is enough to prove the result for one orbit. We may also replace $\phi$ with ${\rm{ad}}_{b}^{k}\phi$ without changing $\Lambda$, so we can assume that $\phi$ and the basis ${\mathcal{S}}$ are as described in Proposition 4.1. Let $s\in\mathcal{S}$ and let $A=\langle s\rangle$. Then $k-1\leq[{\rm{ad}}_{b}^{k}(A)]_{b}^{\mathcal{S}}\leq k+1$ for all $k\in\mathbb{Z}$. We define a map $\mathcal{OF}_{N}\to\mathbb{Z}$ as follows: for each conjugacy class $[B]$ of a free factor we pick a closest point $[\phi]^{j}([A])$ in the $[\phi]$-orbit of $[A]$. This assignment $[B]\mapsto j$ is Lipschitz [BF14], and by composing it with the natural projection $\mathcal{AF}_{N}\to\mathcal{OF}_{N}$ we obtain a Lipschitz map $R\colon\mathcal{AF}_{N}\to\mathbb{Z}$. It follows from Theorem 3.6 that the map $[\,\cdot\,]_{b}^{\mathcal{S}}\times R\colon\mathcal{AF}_{N}\to\mathbb{Z}^{2}$ is also Lipschitz, and we claim that $B\mapsto\phi^{R(B)}{\rm{ad}}_{b}^{[B]_{b}}(A)$ gives a Lipshitz retraction of $\mathcal{AF}_{N}$ onto the orbit of $\Lambda=\langle\phi,{\rm{ad}}_{b}\rangle$. To see this, note that $R(\phi^{j}{\rm{ad}}_{b}^{k}(A))=j$ for all $j,k\in\mathbb{Z}$, and
$k-1-M_{A}\leq[\phi^{j}{\rm{ad}}_{b}^{k}(A)]_{b}\leq k+1+M_{A},$
where $M_{A}$ is the constant provided by Proposition 4.1. ∎
Corollary 4.3.
For $N\geq 2$, the ${\rm{Aut}}$ free factor complex $\mathcal{AF}_{N}$ is not Gromov-hyperbolic.
Proof.
This follows immediately from the theorem and the fact that for every $N\geq 2$ there is a compact surface with one boundary component that has euler characteristic $1-N$ and admits pseudo-Anosov diffeomorphisms. For $N$ even, this is an orientable surface of genus $N/2$, while for $N$ odd we must take the connected sum of an orientable surface of genus $(N-1)/2$ with a projective plane [Pen88]. ∎
5 Problem list
We end the paper with a list of open problems.
Our results extend the list of automorphisms that are known to have undistorted orbits in $\mathcal{AF}_{N}$: Bestvina and Feighn [BF14] proved this is true for fully irredicible automorphisms, and we have added inner automorphisms ${\rm{ad}}_{b}$ given by filling elements.
Question 1.
Which cyclic subgroups of ${\rm{Aut}}(F_{N})$ have undistorted orbits in $\mathcal{AF}_{N}$? Is every orbit of a cyclic subgroup either finite or undistorted?
One can ask about other abelian subgroups of ${\rm{Aut}}(F_{N})$, or more ambitiously, about more general quasi-flats:
Question 2.
Which abelian subgroups of ${\rm{Aut}}(F_{N})$ have undistorted orbits in $\mathcal{AF}_{N}$?
Question 3.
Are there any quasi-flats of rank 3, i.e. (possibly non-equivariant) quasi-isometric embeddings of $\mathbb{Z}^{3}$ into $\mathcal{AF}_{N}$?
We find it unlikely that there should be an action of $\mathbb{Z}^{3}$ with quasi-isometric orbits. The image of any $\mathbb{Z}^{3}\cong\Lambda<{\rm{Aut}}(F_{N})$ in ${\rm{Out}}(F_{N})$ contains a copy of $\mathbb{Z}^{2}$; as we saw in the introduction, this implies that $\Lambda$ has a finite orbit in $\mathcal{OF}_{N}$, and therefore a subgroup of finite index in $\Lambda$ fixes some $[A]\in\mathcal{OF}_{N}$. So if there were a free action of $\mathbb{Z}^{3}$ on $\mathcal{AF}_{N}$ with quasi-isometrically embedded orbits, then there would be such an orbit in the fibre over a point in $\mathcal{OF}_{N}$. It does seem possible, however, that there are $\mathbb{Z}^{2}$ subgroups with undistorted orbits in such a fibre: candidates can be obtained by replacing the pseudo-Anosov $\phi$ in Theroem C with a partial pseudo-Anosov supported on a subsurface containing the boundary of $\Sigma$. In this case $\langle\phi,{\rm{ad}}_{b}\rangle\cong\mathbb{Z}^{2}$ does not fix an obvious free factor, but does fix the conjugacy class of a free factor. It is not clear whether the orbits of such subgroups are undistorted or not.
The quasi-flats in $\mathcal{AF}_{N}$ constructed in this paper come in families where each pair of quasi-flats coarsely share a common line: if $b$ is the boundary word for a surface $\Sigma$, then independent psuedo-Anosov homeomorphisms of $\Sigma$ will give distinct quasi-flats, each pair of which will have bounded neighbourhoods whose intersection is a quasi-line that is a bounded neighbourhood of an $\langle{\rm{ad}}_{b}\rangle$-orbit. This observation allows us to construct many more non-equivariant quasi-isometric embeddings $\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}_{N}$ by piecing together half-flats and parallel strips from the various equivariant flats coarsely containing $\langle{\rm{ad}}_{b}\rangle$-orbits. This pattern of intersections shows that the families of quasi-flats that we currently have cannot be used to define a relative hyperbolic structure on $\mathcal{AF}_{N}$. (We refer to [Sis12] for the definition of such a structure.) But our lack of knowledge about quasi-flats in general means that we cannot answer the following question.
Question 4.
Is $\mathcal{AF}_{N}$ relatively hyperbolic? If not, is it thick in the sense of [BDM09]?
We have seen that if a compact surface $\Sigma$ has one boundary component then every isomorphism $\pi_{1}(\Sigma)\cong F_{N}$ gives rise to a family of quasi-flats that coarsely intersect in a quasi-line. If such families account for all quasi-flats in $\mathcal{AF}_{N}$, then, in the spirit of Dowdall and Taylor’s co-surface graph [DT17], one might hope to obtain a hyperbolic space by coning them off.
Definition 5.1 (An Aut version of the co-surface graph).
Let $\mathcal{AC}_{N}$ be the graph obtained from $\mathcal{AF}_{N}$ by adding a new edge between two points $A$ and $B$ if there exists a surface $\Sigma$ with one boundary component such that both $A$ and $B$ are tethered subsurface subgroups of $\Sigma$ (subsurfaces with an embedded arc to the basepoint, which is on the boundary).
Dowdall and Taylor proved that the co-surface graph of ${\rm{Out}}(F_{N})$ is hyperbolic. Following their lead, we can ask the same question about the Aut version.
Question 5.
Is $\mathcal{AC}_{N}$ Gromov-hyperbolic?

### 12 Mar 2024 v2

@@
 Abstract
-Let $\Phi$ be a pseudo-Anosov diffeomorphism of a compact (possibily non-orientable) surface $\Sigma$ with one boundary component.
+Let $\Phi$ be a pseudo-Anosov diffeomorphism of a compact (possibly non-orientable) surface $\Sigma$ with one boundary component.
 We show that if $b\in\pi_{1}(\Sigma)$ is the boundary word, $\phi\in{\rm{Aut}}(\pi_{1}(\Sigma))$ is a representative of $\Phi$ fixing $b$, and ${\rm{ad}}_{b}$ denotes conjugation by $b$, then the orbits of $\langle\phi,{\rm{ad}}_{b}\rangle\cong\mathbb{Z}^{2}$ in the graph of free factors of $\pi_{1}(\Sigma)$ are quasi-isometrically embedded.
@@
 The natural map ${\rm{Out}}(F_{N})\to{\rm{Isom}}(\mathcal{OF}_{N})$ is an isomorphism [BB23], so Theorem A describes all possible actions of cyclic groups on $\mathcal{OF}_{N}$.
-As the graph is hyperbolic, we also know that $\mathbb{Z}^{2}$ cannot with act quasi-isometric orbits.
+As the graph is hyperbolic, we also know that $\mathbb{Z}^{2}$ cannot act with quasi-isometric orbits.
 In fact, results in the literature [BB23, HM20] tell us something much stronger:
@@
 Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$.
-Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
+Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $\lambda\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
 Corollary D.
@@
 A crucial property of the boundary element $b\in F_{N}$ is that it is filling in the sense that it is not contained in any proper free factor of $F_{N}$.
-Given any filling element $b$, we analyse the action of ${\rm{ad}}_{b}$ on $\mathcal{AF}_{N}$ with the aim or showing that the orbits of $\langle{\rm{ad}}_{b}\rangle$ give quasi-isometric embeddings of $\mathbb{Z}$.
+Given any filling element $b$, we analyse the action of ${\rm{ad}}_{b}$ on $\mathcal{AF}_{N}$ with the aim of showing that the orbits of $\langle{\rm{ad}}_{b}\rangle$ give quasi-isometric embeddings of $\mathbb{Z}$.
 The key tool in this analysis is the notion of $b$-reduced decomposition for elements $w\in F_{N}$, which we introduce in Section 3.
@@
 This paper is dedicated to Slava Grigorchuk on his 70th birthday, with thanks for his many highly-original contributions to group theory.
+We thank the anonymous referee for their helpful comments.
 The first author acknowledges the support of the National Science Foundation under the grant number DMS-2304774.
@@
 This is the graph with vertex set $\mathcal{S}\cup\mathcal{S}^{-1}$ that has an edge from $x$ to $y^{-1}$ if $xy$ is a subword of the cyclic reduction $w_{0}$ or if $x$ is the last letter of $w_{0}$ and $y$ is the first letter of $w_{0}$.
+Importantly for us, the cyclic Whitehead graph keeps track of the turns crossed by the axis of the element $w$ in the Cayley tree determined by $\mathcal{S}$.
 A vertex $v$ of a graph is a cut vertex if the full subgraph spanned by vertices not equal to $v$ is disconnected.
@@
 Fix a basis $\mathcal{S}$ of $F_{N}$.
-Given a cyclically reduced word $b$ and a reduced word $w$, representing elements of $F_{N}$, the $b$-reduced decomposition of $w$ (with respect to $\mathcal{S}$) is the decomposition
+Given a cyclically reduced word $b$ and a reduced word $w$ representing elements of $F_{N}$, the $b$-reduced decomposition of $w$ (with respect to $\mathcal{S}$) is the decomposition
 $w\equiv b^{k}w_{b}b^{-k}$
@@
 Proof.
+As $\mathcal{S}$ minimizes $|b|_{\mathcal{S}}$, in particular $b$ is cyclically reduced.
 Let $T$ be the Cayley tree for $F_{N}$ with respect to the basis $\mathcal{S}$.
@@
 The intersection of $p$ with $X_{b}$ is of length at most $|b|_{\mathcal{S}}-1$, as the reduced word representing $a$ starts with the word labelling $p$ and ends with its inverse and $[a]_{b}=0$.
-Furthermore, the intersection of $X_{a}$ with $X_{b}$ can have length at most $|b|_{\mathcal{S}}$: to see this, note that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ does not contain a cut vertex by Proposition 2.2, whereas ${\mathrm{Wh}}_{\mathcal{S}}(a)$ does by the Cut Vertex Lemma, so $X_{a}$ cannot contain all the turns (pairs of consecutive of labels from $\mathcal{S}\cup\mathcal{S}^{-1}$) that appear along $X_{b}$.
+Furthermore, the intersection of $X_{a}$ with $X_{b}$ can have length at most $|b|_{\mathcal{S}}$: to see this, note that the cyclic Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ does not contain a cut vertex by Proposition 2.2, whereas ${\mathrm{Wh}}_{\mathcal{S}}(a)$ does by the Cut Vertex Lemma, so $X_{a}$ cannot contain all the turns (pairs of consecutive of labels from $\mathcal{S}\cup\mathcal{S}^{-1}$) that appear along $X_{b}$.
 Let $y$ by the furthest point along $X_{b}$ in the $b^{-1}$ direction that is contained in $p\cup X_{a}$.
@@
 Proposition 3.5.
-Suppose $b\in F_{N}$ is cyclically reduced with respect to ${\mathcal{S}}$ and that the Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ is connected with no cut vertex.
+Suppose $b\in F_{N}$ is cyclically reduced with respect to ${\mathcal{S}}$ and that the cyclic Whitehead graph ${\mathrm{Wh}}_{\mathcal{S}}(b)$ is connected with no cut vertex.
 Let $T$ be the Cayley tree with respect to $\mathcal{S}$ and let $X_{b}$ be the axis of $b$ in $T$.
@@
 As in Proposition 3.5(1), we know that the projection $\pi_{X_{b}}(X_{a_{i}})$ of each axis to $X_{b}$ has length at most $|b|_{\mathcal{S}}$.
-We claim that furthermore, if the projections are disjoint then the arc connecting $\pi_{X_{b}}(X_{a_{1}})$ to $\pi_{X_{b}}(X_{a_{2}})$ is length at most $|b|_{\mathcal{S}}$.
+We claim that, furthermore, if the projections are disjoint then the arc connecting $\pi_{X_{b}}(X_{a_{1}})$ to $\pi_{X_{b}}(X_{a_{2}})$ is length at most $|b|_{\mathcal{S}}$.
 This claim clearly gives a uniform bound on the difference between $[a_{1}]_{b}$ and $[a_{2}]_{b}$, and with a similar case-by-case analysis as in Proposition 3.5 one can check that $[a_{1}]_{b}$ and $[a_{2}]_{b}$ differ by at most $2$.
@@
 Let $f:T^{\mathcal{S}}\to T^{\mathcal{T}}$ be the map that sends each edge of $T^{\mathcal{S}}$ to the geodesic in $T^{\mathcal{T}}$ with the same endpoints.
-If $p=[1,g]_{\mathcal{S}}$ is a path in $T^{\mathcal{S}}$ from 1 to the axis of an element $a\in F_{N}$, then the bounded backtracking property (see [Coo87, GJLL98]) implies that $g$ lies at distance at most $C$ from the axis of $a$ in $T^{\mathcal{T}}$, where $C$ is the bounded backtracking constant.
-Therefore, as $|b|_{\mathcal{S}}=|b|_{\mathcal{T}}$, the proposition reduces to showing that for any path $p$ based at 1, the reduction of $f(p)$ spends approximately the same amount of time travelling along the axis of $b$ in $T^{\mathcal{T}}$ as the path $p$ does along the axis of $b$ in $T^{\mathcal{S}}$ (in the same direction).
-This is reasonably standard: decompose such a path $p$ as $p=p_{1}p_{2}p_{3}$, where $p_{1}=[1,b^{k}]_{\mathcal{S}}$ for some $k$, the path $p_{2}$ is a segment of length $\leq|b|_{\mathcal{S}}$, and $p_{3}$ is a path disjoint from the axis $X^{\mathcal{S}}_{b}\subset T^{\mathcal{S}}$ except for its initial endpoint.
+Let $1_{\mathcal{S}}$ and $1_{\mathcal{T}}$ be the basepoints of the respective Cayley trees and use $g_{\mathcal{S}}=g.1_{\mathcal{S}}$ and $g_{\mathcal{T}}=g.1_{\mathcal{T}}$ to denote the vertices in the trees given by $g\in F_{N}$.
+If $p=[1_{\mathcal{S}},g_{\mathcal{S}}]$ is a path in $T^{\mathcal{S}}$ from $1_{\mathcal{S}}$ to the axis of an element $a\in F_{N}$, then the bounded backtracking property (see [Coo87, GJLL98]) implies that $g_{\mathcal{T}}=f(g_{\mathcal{S}})$ lies at distance at most $C$ from the axis of $a$ in $T^{\mathcal{T}}$, where $C$ is the bounded backtracking constant.
+Therefore, as $|b|_{\mathcal{S}}=|b|_{\mathcal{T}}$, the proposition reduces to showing that for any path $p$ based at $1_{\mathcal{S}}$, the reduction of $f(p)$ spends approximately the same amount of time travelling along the axis of $b$ in $T^{\mathcal{T}}$ as the path $p$ does along the axis of $b$ in $T^{\mathcal{S}}$ (in the same direction).
+This is reasonably standard: decompose such a path $p$ as $p=p_{1}p_{2}p_{3}$, where $p_{1}=[1_{\mathcal{S}},b_{\mathcal{S}}^{k}]$ for some $k$, the path $p_{2}$ is a segment of length $\leq|b|_{\mathcal{S}}$, and $p_{3}$ is a path disjoint from the axis $X^{\mathcal{S}}_{b}\subset T^{\mathcal{S}}$ except for its initial endpoint.
 As $f$ is bi-Lipschitz on the vertex set, $p_{3}$ can further be decomposed as $p_{3}^{\prime}p_{3}^{\prime\prime}$, where $f(p_{3}^{\prime\prime})$ is disjoint from $X^{\mathcal{T}}_{b}$ and $p_{3}^{\prime}$ is of uniformly bounded length.
-Then the intersection of the reduction of $f(p)$ with $X^{\mathcal{T}}_{b}$ in $T_{\mathcal{T}}$ is $[1,b^{k}]_{\mathcal{T}}$, up to cancellation/concatenation with a subpath of $f(p_{2}p_{3}^{\prime})$, which is of uniformly bounded length.
+Then the intersection of the reduction of $f(p)$ with $X^{\mathcal{T}}_{b}$ in $T_{\mathcal{T}}$ is $[1_{\mathcal{T}},b_{\mathcal{T}}^{k}]$, up to cancellation/concatenation with a subpath of $f(p_{2}p_{3}^{\prime})$, which is of uniformly bounded length.
 ∎
@@
 Proof.
-In Sections 1-3 of [BH92], it is shown that an automorphism $\phi_{0}$ as above has a stable train track representative $f\colon G\to G$ with exactly one indivisible Nielsen path (iNP) $\rho$ which, as the automorphism is geometric, forms a loop representing $b$ based at a point $x\in G$ that is fixed by $f$.
+In Sections 1-3 of [BH92], it is shown that an automorphism $\phi_{0}$ as above has a stable train track representative $f\colon G\to G$ on a graph $G$ with exactly one indivisible Nielsen path (iNP) $\rho$ which, as the automorphism is geometric, forms a loop representing $b$ based at a point $x\in G$ that is fixed by $f$.
 Furthermore, $\rho$ crosses every edge in $G$ exactly twice (see [BH92, Lemma 3.9] and the discussion in [BH92, Section 4]).
-Identifying $\pi_{1}(G,x)$ with $F_{N}$, we define $\phi$ to be the automorphism induced by $f_{*}\colon\pi_{1}(G,x)\to\pi_{1}(G,x)$ and take $\mathcal{S}$ to be a basis of $F_{N}$ determined by a choice of maximal tree $T$ in $G$ together with an orientation of the edges in $G\smallsetminus T$.
+We identify $\pi_{1}(G,x)$ with $F_{N}$ via a choice of maximal tree $T$ in $G$ together with an orientation of the edges in $G\smallsetminus T$.
 Each edge-loop based at $x$ determines a word in the letters ${\mathcal{S}}^{\pm 1}$ whose length is the number of edges of $G\smallsetminus T$ crossed by the loop.
 In particular, $|b|_{\mathcal{S}}=2N$.
-Since $b$ is the boundary for a surface of Euler characteristic $1-N$, this tells us that $\mathcal{S}$ minimizes the length of $b$.
+Since $b$ is the boundary of a surface of Euler characteristic $1-N$, this tells us that $\mathcal{S}$ minimizes the length of $b$.
+We define $\phi$ to be the automorphism induced by $f_{*}\colon\pi_{1}(G,x)\to\pi_{1}(G,x)$.
 We now prove the second part with $r\geq 0$.
@@
 As $b^{k}\tilde{x}$ is also fixed under $\tilde{f}$, this implies that $k-1\leq[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}\leq k+1$ for all $k\in\mathbb{Z}$ and $r\geq 0$.
-As the map $\mathcal{AF}_{N}\to\mathbb{Z}$ induced by $[A]\mapsto[A]_{b}$ is 2-Lipschitz, this implies that for all $A\in\mathcal{AF}_{N}$ there exists $C_{A}$ such that $|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq C_{A}$ for all $r\geq 0$ and $k\in\mathbb{Z}$.
+As the map $\mathcal{AF}_{N}\to\mathbb{Z}$ induced by $[A]\mapsto[A]_{b}$ is 2-Lipschitz, we have
+$\begin{split}|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\leq{}&|[\phi^{r}{\rm{ad}}_{b}^{k}(A)]_{b}-[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}|+|[\phi^{r}{\rm{ad}}_{b}^{k}(l)]_{b}-[{\rm{ad}}_{b}^{k}(l)]_{b}|\\ &+|[{\rm{ad}}_{b}^{k}(l)]_{b}-[{\rm{ad}}_{b}^{k}(A)]_{b}|\end{split}$
+[…]
+for all $r\geq 0$ and $k\in\mathbb{Z}$.
 Note that any possible choices for representatives of $[\phi_{0}]$ fixing $b$ differ by a power of ${\rm{ad}}_{b}$.
@@
 Note that although you might be working with a different basis $\mathcal{S}^{\prime}$ on which $b$ is minimal length, the projections $\mathcal{AF}_{N}\to\mathbb{Z}$ are coarsely equivalent by Proposition 3.7.
-Furthermore, this is the unique representative of $[\phi^{\prime}]^{-1}$ fixing $b$ such that $b^{\infty}$ and $b^{-\infty}$ are nonattracting under $\partial\psi$.
+Furthermore, this is the unique representative of $[\phi_{0}]^{-1}$ fixing $b$ such that $b^{\infty}$ and $b^{-\infty}$ are nonattracting under $\partial\psi$.
 This implies that $\psi=\phi^{-1}$, so we can finish part (2) by taking $M_{A}:=\max\{C_{A},C_{A}^{\prime}\}$.
@@
 Let $\Sigma$ be a compact surface with one boundary component, fix an identification $\pi_{1}\Sigma=F_{N}$, let $b\in F_{N}$ be the homotopy class of the boundary loop, let $\phi\in{\rm{Aut}}(F_{N})$ be the automorphism induced by a pseudo-Anosov element of $\text{{\rm{Mod}}}(\Sigma)$, let $\Lambda=\langle\phi,\,{\rm{ad}}_{b}\rangle<{\rm{Aut}}(F_{N})$ and note that $\Lambda\cong\mathbb{Z}^{2}$.
-Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $A\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
+Then, for every $A\in\mathcal{AF}_{N}$, the orbit map $\lambda\mapsto\lambda(A)$, defines a quasi-isometric embedding $\Lambda\cong\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}$.
 Proof.
@@
 We define a map $\mathcal{OF}_{N}\to\mathbb{Z}$ as follows: for each conjugacy class $[B]$ of a free factor we pick a closest point $[\phi]^{j}([A])$ in the $[\phi]$-orbit of $[A]$.
-This assignment $[B]\mapsto j$ is Lipschitz [BF14], and by composing it with the natural projection $\mathcal{AF}_{N}\to\mathcal{OF}_{N}$ we obtain a Lipschitz map $R\colon\mathcal{AF}_{N}\to\mathbb{Z}$.
+This assignment $[B]\mapsto j$ is Lipschitz [BF14], and by composing it with the natural projection $\mathcal{AF}_{N}\to\mathcal{OF}_{N}$ we obtain a Lipschitz map $R\colon\mathcal{AF}_{N}\to\mathbb{Z}$ which is invariant under $\text{Inn}(F_{N})$.
 It follows from Theorem 3.6 that the map $[\,\cdot\,]_{b}^{\mathcal{S}}\times R\colon\mathcal{AF}_{N}\to\mathbb{Z}^{2}$ is also Lipschitz, and we claim that $B\mapsto\phi^{R(B)}{\rm{ad}}_{b}^{[B]_{b}}(A)$ gives a Lipshitz retraction of $\mathcal{AF}_{N}$ onto the orbit of $\Lambda=\langle\phi,{\rm{ad}}_{b}\rangle$.
@@
 We end the paper with a list of open problems.
-Our results extend the list of automorphisms that are known to have undistorted orbits in $\mathcal{AF}_{N}$: Bestvina and Feighn [BF14] proved this is true for fully irredicible automorphisms, and we have added inner automorphisms ${\rm{ad}}_{b}$ given by filling elements.
+Our results extend the list of automorphisms that are known to have undistorted orbits in $\mathcal{AF}_{N}$: Bestvina and Feighn [BF14] proved this is true for fully irreducible automorphisms, and we have added inner automorphisms ${\rm{ad}}_{b}$ given by filling elements.
 Question 1.
@@
 It is not clear whether the orbits of such subgroups are undistorted or not.
-The quasi-flats in $\mathcal{AF}_{N}$ constructed in this paper come in families where each pair of quasi-flats coarsely share a common line: if $b$ is the boundary word for a surface $\Sigma$, then independent psuedo-Anosov homeomorphisms of $\Sigma$ will give distinct quasi-flats, each pair of which will have bounded neighbourhoods whose intersection is a quasi-line that is a bounded neighbourhood of an $\langle{\rm{ad}}_{b}\rangle$-orbit.
+The quasi-flats in $\mathcal{AF}_{N}$ constructed in this paper come in families where each pair of quasi-flats coarsely share a common line: if $b$ is the boundary word for a surface $\Sigma$, then independent pseudo-Anosov homeomorphisms of $\Sigma$ will give distinct quasi-flats, each pair of which will have bounded neighbourhoods whose intersection is a quasi-line that is a bounded neighbourhood of an $\langle{\rm{ad}}_{b}\rangle$-orbit.
 This observation allows us to construct many more non-equivariant quasi-isometric embeddings $\mathbb{Z}^{2}\hookrightarrow\mathcal{AF}_{N}$ by piecing together half-flats and parallel strips from the various equivariant flats coarsely containing $\langle{\rm{ad}}_{b}\rangle$-orbits.